OpenStudy (anonymous):

I am really struggling re gradients of a straight line, please someone help! Q. Write down the gradient, m, for each of these straight lines: 1) y= 3-3x ii) y= 2+x/3 iii) 3x+5y-8=0 iv) 2y+3x=0 v) 5x-4y-12=0 vi) 7x+6y-18=0 vii) 5x-2y/3-1=0 If someone could explain to me how this is done I would really appreciate it, thanks in advance.

OpenStudy (anonymous):

Y=MX+C in which the M stands for the Gradient and the C for the Y intercept. Rearrange all of the above questions into the structure Y=MX + C by isolating Y. The coefficient of X is the Gradient E.g Question one is already arranged , coefficient of X is -3. Thus -3 is the answer iii) 3x+5y-8=0 Rearranges to 5Y= 8-3x thus Y=8/5 - (3/5 x) coefficient of X is -3/5 and thus -3/5 is the answer

OpenStudy (anonymous):

Yes, thanks a ton :D